OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and perpendicular to y = 1 fifthx - 10. y = 5x + 10 y = 5x y = -5x - 10 y = -5x

OpenStudy (anonymous):

this one?

OpenStudy (anonymous):

y = 1 fifthx - 10 y = (1/5) x - 10 so we need to be perpendicular to this line, so the steps were to take the slope form this line, y = (1/5) x - 10, turn it into a fraction (its already a fraction so skip this step), flip the fraction upside down. and times it by -1. This will give us the slope of the perpendicular line. can you do this, we did it before on the last one?

OpenStudy (anonymous):

I think so. So 1 / 5 flipped is 5 /1, right? Then its opposite would be - 5 /1? But you could also just write it as -5? Right :) @DemolisionWolf

OpenStudy (anonymous):

excellent work, -5 is correct. so, we now have m=-5 we create our new line equation the same way as last time, we take the standard form of y=mx+b plug in m, y=-5x+b plug in the point our new line has to pass through, (1, -5) -5 = -5(1) + b now solve for b..

OpenStudy (anonymous):

0 = b?

OpenStudy (anonymous):

So it would be y = -5x?

OpenStudy (anonymous):

My os account is being dumb, so not to be annoying, but I'm gonna mention you just to make sure my messages go through :P @DemolisionWolf

OpenStudy (anonymous):

y=-5x is correct ^_^

OpenStudy (anonymous):

my account is being dumb too

OpenStudy (anonymous):

thanks a ton!!!! Can you see anyone's smartscores?

OpenStudy (anonymous):

Can you make sure I did this right? Write the equation of the line in slope-intercept form with a slope of 2 and a y-intercept of -3. I have y = 2x + -3

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

I can't see anyon'es smart score, nor do I get my notifications

OpenStudy (anonymous):

y = 2x + -3 correct

OpenStudy (anonymous):

Ditto, I think theres a problem with open study for everyone.

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

yep ^_^